the edge length of the fcc unit cell is 508pm.if the lradious of ht ecation is 110pm the radious of the anion is

the edge  length of the fcc unit cell is 508pm.if the lradious of ht ecation is 110pm the radious of the anion is
 
 

Grade:12th pass

1 Answers

Arun
5 years ago

Centered unit cell -

In centered unit cell, constituent particles are present in position other than corners in addition to corners.

- wherein

Three types: face centered, body centered, end centered.

 

 For fcc unit cell, 2 (radius of cation+ radius of anion) = edge length 

Therefore (110+ra) = (110+r_{a})=\frac{508}{2}\Rightarrow r_{a}=254-110= 144pm